OpenStudy (anonymous):

x=4 asolution for 3x-1=11

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

12-1=11

OpenStudy (anonymous):

yes 3(4)-1=11

OpenStudy (anonymous):

how did you get 12

OpenStudy (anonymous):

3x-1=11 x=4 3(4)-1=11 12-1=11

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

vote for me

OpenStudy (anonymous):

click on the "good answer" box